5.88-10-10 3 Constants 1 Periodic Table Part B A concert loudspeaker suspended high off the ground emits 25 0 W of sound power. A small microphone with a 0 800 m2 area is 52 0 m from the speaker How much sound energy impinges on the microphone each secondr Express your answer with the appropriate units. Enter your answer with a different unik type Revlew a lst of acceptable units ANSWER 2 Deduction-3% 5.88-10 dB x 10 your answer with a different unit type. Review a list of acceptable unts 5.88.101J ANSWER 3 Dedeton-3% Submit 6.88-10 10 J
Answer #1

(A) Sound intensity: I at a distance remote from the source follows an inverse square law.

I = P/A = P/4πr² = 25/4π52² = 7.36*10^-4 W/m²

(B) Energy incident per m^2 per second is intensity I

Energy incident per 0.8x10^-4m^2 per second is

I x 0.8x10^-4 = 7.36x10^-4x0.8x10^-4 = 5.89*10^-8 J